# Fisher-Rao Gradient Flow: Geodesic Convexity and Functional Inequalities

### v1

@@ 5.3. General results using the reverse $\chi^{2}$ divergence as the dual
(5.13) $\begin{split}&\int\Bigl{(}f^{\prime}\Bigl{(}\frac{1}{\mu}\Bigr{)}-f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\Bigr{)}\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\rho\mathrm{d}\theta\\ \geq&\mu(1-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}f(\frac{1}{\mu})=\mu^{2}h(\frac{1}{\mu})\quad(\textrm{using }\mu\geq 1+\delta_{f}\textrm{ and }h\textrm{non-increasing},h\geq 0)\\ \geq&h(\frac{1}{1+\delta_{f}})(\mu^{2}-1)=h(\frac{1}{1+\delta_{f}})\chi^{2}[{\rho^{*}}\|\rho].\end{split}$
According to the definition of $\delta_{f}$ in (5.9), we have
(5.14) $h(\frac{1}{1+\delta_{f}})=\int^{1}_{\frac{1}{1+\delta_{f}}}f^{\prime\prime}(x)(1-x)\mathrm{d}x\geq\frac{f^{\prime\prime}(1)}{2}\int^{1}_{\frac{1}{1+\delta_{f}}}(1-x)\mathrm{d}x=\frac{\delta_{f}^{2}f^{\prime\prime}(1)}{4(1+\delta_{f})^{2}}.$
@@
Combining the two scenarios, namely $\mu\leq 1+\delta_{f}$ in (5.11) and $\mu\geq 1+\delta_{f}$ in (LABEL:eq:generalized-gradient-dominant-condition-2b), we conclude that (5.4b) holds for
$\alpha_{f}=\frac{1}{2}\min\{1,\alpha^{\prime}_{f},\frac{\delta_{f}^{2}f^{\prime\prime}(1)}{4(1+\delta_{f})^{2}}\}>0.$
This completes the proof. ∎
5.4. Tighter results using the *-conjugate $f$-divergence as the dual
@@
In the last subsection, we obtain general convergence results by choosing $\bar{f}$ as the reverse $\chi^{2}$ divergence. The constant $\alpha_{f}$ depends on $f$ only. However, it depends on some implicit constant such as $\delta_{f},\alpha_{f}^{\prime}$ that are dependent on $f$. The goal of this subsection is to improve the estimate for specific cases and obtain more explicit constants.
@@
We take ${\bar{f}}(x)=xf\bigl{(}\frac{1}{x}\bigr{)}$, known as as the *-conjugate function of $f$ [67]. For this choice, the $\bar{f}$-divergence has the following dual form
(5.15) $D_{{\bar{f}}}[\rho\|{\rho^{*}}]=\int{\rho^{*}}{\bar{f}}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\,\mathrm{d}\theta=\int\rho f\Bigl{(}\frac{{\rho^{*}}}{\rho}\Bigr{)}\,\mathrm{d}\theta.$
@@
Along the Fisher-Rao gradient flow (2.3) with respect to $D_{f}$, we have
@@
In the following theorem, we show the dual gradient dominance condition holds with an explicit constant, when $f^{\prime\prime}$ is a polynomial.
@@
Given any convex function $f$ specified by $f^{\prime\prime}(x)=x^{p}$ with $f(1)=0$, when $p\leq-2$ or $-1\leq p$, we have
(5.16) $-\int\Bigl{(}-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}\Bigl{(}f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr{)}\mathrm{d}\theta\geq D_{\bar{f}}[\rho\|{\rho^{*}}]+D_{f}[\rho\|{\rho^{*}}],$
and when $-2<p<-1$, we have
(5.17) $-\int\Bigl{(}-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}\Bigl{(}f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr{)}\mathrm{d}\theta\geq\frac{1}{2}\Bigl{(}D_{\bar{f}}[\rho\|{\rho^{*}}]+D_{f}[\rho\|{\rho^{*}}]\Bigr{)}.$
@@
Remark 5.7.
As a consequence of Theorem 5.6, we get a convergence result for the Fisher-Rao gradient flow associated with the KL-divergence. More precisely, for the Fisher-Rao gradient flow of the KL-divergence
(5.18) $\frac{\partial\rho_{t}}{\partial t}=-\rho_{t}\Bigl{(}\bigl{(}\log\rho_{t}-\log{\rho^{*}}\bigr{)}-\mathbb{E}_{\rho_{t}}[\log\rho_{t}-\log{\rho^{*}}]\Bigr{)},$
if we choose an initial density $\rho_{0}$ satisfying that
${\rm KL}[\rho_{0}\|{\rho^{*}}]+{\rm KL}[{\rho^{*}}\|\rho_{0}]<\infty,$
then the Fisher-Rao gradient flow (assuming it is well-posed) converges exponentially fast with
@@
$\lozenge$
Proof of Theorem 5.6.
First, we observe that both sides of these dual gradient dominance inequalities (5.16) and (5.17) remain the same under the transformation $f(x)\rightarrow f(x)+c(x-1)$ for any $c$. Therefore, for each $p$, it suffices to prove that (5.16) and (5.17) hold for one representative element under the transformation.
When $p=-1$, we consider $f(x)=x\log x$, the corresponding $f$-divergence is the KL-divergence. We have
(5.19) $-\int\Bigl{(}-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}\Bigl{(}f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr{)}\mathrm{d}\theta$
@@
Hence, (5.16) holds.
When $p=-2$, we consider $f(x)=-\log x$, the corresponding $f$-divergence is the reverse KL-divergence. We have
(5.20) $-\int\Bigl{(}-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}\Bigl{(}f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr{)}\mathrm{d}\theta$
@@
Hence, (5.16) holds.
When $p\neq-1$ and $p\neq-2$, we consider $f(x)=\frac{x^{p+2}-1}{(p+2)(p+1)}$. The corresponding $f$-divergences include $\chi^{2}$ distance (when $p=0$). We have
(5.21) $-\int\Bigl{(}-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}\Bigl{(}f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr{)}\mathrm{d}\theta$
@@
When $p<-2$ or $-1<p$, applying Jensen’s inequality to the convex function $x^{p+2}~{}(x>0)$, we have
$\int\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}\mathrm{d}\theta=\mathbb{E}_{{\rho^{*}}}\Bigl{[}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}^{p+2}\Bigr{]}\geq 1\quad\textrm{ and }\quad\int\frac{{\rho^{*}}^{p+2}}{\rho^{p+1}}\mathrm{d}\theta=\mathbb{E}_{\rho}\Bigl{[}\bigl{(}\frac{{\rho^{*}}}{\rho}\bigr{)}^{p+2}\Bigr{]}\geq 1.$
Therefore, (5.16) holds using the following estimate:
(5.22) $\frac{1}{(p+2)(p+1)}$ $\Bigl{(}\int\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}\mathrm{d}\theta\int\frac{{\rho^{*}}^{p+2}}{\rho^{p+1}}\mathrm{d}\theta-1\Bigr{)}$
@@
When $-2<p<-1$, applying Jensen’s inequality to the concave function $x^{p+2}~{}(x>0)$, we have
$\int\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}\mathrm{d}\theta=\mathbb{E}_{{\rho^{*}}}\Bigl{[}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}^{p+2}\Bigr{]}\leq 1\quad\textrm{ and }\quad\int\frac{{\rho^{*}}^{p+2}}{\rho^{p+1}}\mathrm{d}\theta=\mathbb{E}_{\rho}\Bigl{[}\bigl{(}\frac{{\rho^{*}}}{\rho}\bigr{)}^{p+2}\Bigr{]}\leq 1.$
Note that $(p+1)(p+2)<0$. We get that (5.17) using the following estimate:
(5.23) $\frac{1}{(p+2)(p+1)}$ $\Bigl{(}\int\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}\mathrm{d}\theta\int\frac{{\rho^{*}}^{p+2}}{\rho^{p+1}}\mathrm{d}\theta-1\Bigr{)}$
@@
This concludes the proof. ∎
6. Conclusion
@@
Furthermore, we have introduced the dual gradient dominance condition and demonstrated its validity for a broad class of $f$-divergences, notably including the KL divergence. The constants in this condition depend solely on $f$, independent of the target distribution. This characteristic sharply contrasts with the log-Sobolev inequality developed in the Wasserstein geometry, where constants depend on the target distribution.
@@
Future research directions include establishing the well-posedness of the Fisher-Rao gradient flows and exploring the properties of more general, including non-smooth, energy functionals within the Fisher-Rao geometry and the related Wasserstein-Fisher-Rao geometry. Designing efficient numerical approximations for these Fisher-Rao gradient flows, particularly for sampling applications as discussed in [59, 60, 44, 56, 22, 63], remains a promising and open area of research.
Acknowledgments JAC was supported by the Advanced Grant Nonlocal-CPD (Nonlocal PDEs for Complex Particle Dynamics: Phase Transitions, Patterns and Synchronization) of the European Research Council Executive Agency (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 883363). JAC was also partially supported by the EPSRC grants EP/T022132/1 and EP/V051121/1 and by the “Maria de Maeztu” Excellence Unit IMAG, reference CEX2020-001105-M, funded by MCIN/AEI/10.13039/501100011033/. YC was supported by the Courant instructorship. DZH was supported by “The Fundamental Research Funds for the Central Universities, Peking University”. We thank De Huang for useful discussions on log-Sobolev inequalities.
@@
Proof of Lemma 4.5.
In this lemma, our assumptions (conditions (4.6) and (4.19)) are that there exists $10<\delta<$ 1, and $0<\alpha_{\delta}\leq\alpha$, such that
(A.1) $\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}^{2}\geq\alpha_{\delta}(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)},\ \forall\ x\in(1-\delta,1+\delta)\ \textrm{and}\ y>0,$
where the above inequality also holds for any $0<x<1<y$. We also have the constraint
(A.2) $\sum_{i=1}^{3}\rho_{i}=1,\quad\sum_{i=1}^{3}\frac{\rho_{i}}{x_{i}}=1,\quad\rho_{i}>0,\quad x_{i}>0.$
Our goal is to show
(A.3) $\sum_{1\leq i,j\leq 3}\rho_{i}\rho_{j}\Bigl{(}f^{\prime}(x_{i})-f^{\prime}(x_{j})\Bigr{)}^{2}\geq\alpha_{f}\sum_{1\leq i,j\leq 3}\rho_{i}\rho_{j}(\frac{1}{x_{i}}-\frac{1}{x_{j}})\Bigl{(}\frac{f(x_{i})}{x_{i}-1}-\frac{f(x_{j})}{x_{j}-1}\Bigr{)}.$
We assume that $x_{1}\leq x_{2}\leq x_{3}$, then due to the constraint (A.2), we have either $x_{1}=x_{2}=x_{3}=1$ or $x_{1}<1<x_{3}$.
For the first scenario, (A.3) holds with any $\alpha_{f}$. For the second scenario, when $1-\delta<x_{2}<1+\delta$, the relationships between pairs $(x_{1},x_{2})$, $(x_{1},x_{3})$ and $(x_{1},x_{3})$ fall into the conditions described in (A.1). Thus, (A.3) holds with $\alpha_{f}=\alpha_{\delta}$. Now, we only need to verify the following cases:
  1. Case 1
    . $0<x_{1}<1<1+\delta\leq x_{2}\leq x_{3}$,
  2. Case 2
    . $0<x_{1}\leq x_{2}\leq 1-\delta<1<x_{3}$.
For Case 1, we will first show that
(A.4) $\rho_{1}\rho_{3}(\frac{1}{x_{1}}-\frac{1}{x_{3}})\Bigl{(}\frac{f(x_{1})}{1-x_{1}}-\frac{f(x_{3})}{1-x_{3}}\Bigr{)}\geq\delta\rho_{2}\rho_{3}(\frac{1}{x_{2}}-\frac{1}{x_{3}})\Bigl{(}\frac{f(x_{2})}{1-x_{2}}-\frac{f(x_{3})}{1-x_{3}}\Bigr{)}.$
In fact, the convexity of $f(x)$ implies that $\frac{f(x)}{1-x}$ is decreasing. Hence eq. A.4 can be obtained if we can show $\rho_{1}(\frac{1}{x_{1}}-\frac{1}{x_{3}})\geq\delta\rho_{2}(\frac{1}{x_{2}}-\frac{1}{x_{3}}).$ This can be achieved using the following estimate:
Here in the last inequality, we used the fact that $x_{2}\geq 1+\delta$ and $x_{3}>0$.
With (A.4), we can prove that (A.3) holds with $\alpha_{f}=\frac{\alpha_{\delta}\delta}{1+\delta}$. To do so, we denote
(A.5) $F_{i,j}=\rho_{i}\rho_{j}(\frac{1}{x_{i}}-\frac{1}{x_{j}})\Bigl{(}\frac{f(x_{i})}{1-x_{i}}-\frac{f(x_{j})}{1-x_{j}}\Bigr{)},$
then we have
[…]
which shows that (A.3) holds with $\alpha_{f}=\frac{\alpha_{\delta}\delta}{1+\delta}$.
For Case 2, we will first show that
(A.6) $\rho_{1}\rho_{3}(\frac{1}{x_{1}}-\frac{1}{x_{3}})\Bigl{(}\frac{f(x_{1})}{1-x_{1}}-\frac{f(x_{3})}{1-x_{3}}\Bigr{)}\geq\delta\rho_{2}\rho_{1}(\frac{1}{x_{1}}-\frac{1}{x_{2}})\Bigl{(}\frac{f(x_{1})}{1-x_{1}}-\frac{f(x_{2})}{1-x_{2}}\Bigr{)}.$
In fact, the convexity of $f(x)$ implies that $\frac{f(x)}{1-x}$ is decreasing, and hence (A.6) can be obtained if we have $\rho_{3}\geq\delta\rho_{2}$. This can be derived by the following estimate:
Here, in the last inequality, we used the fact that $x_{2}\leq 1-\delta$. Next, we will prove that (A.3) holds with $\alpha_{f}=\frac{\alpha_{\delta}\delta}{1+\delta}$. Using the definition in (A.5), we have
[…]
Combining both Case 1 and Case 2, we complete the proof. ∎
Proof of Lemma 4.4.
Assume that 4.2 holds with $\alpha>0$, namely
(A.7) $\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}^{2}\geq\alpha(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)}$
holds for any $0<x<1<y$. Let $y$ in (A.7) approach $1$, we get
(A.8) $f^{\prime}(x)^{2}\geq\alpha\frac{f(x)}{x}\quad\textrm{ for }\quad x>0.$
(A.9) $f^{\prime\prime}(1)\geq\frac{\alpha}{2}.$
(A.10) $\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x)\leq 2f^{\prime\prime}(1).$
(A.11) $-f^{\prime}(1-\delta)\leq\frac{\delta^{\prime}}{2}f^{\prime\prime}(1).$
Also, the convexity of $f$ implies that $f^{\prime}(x)$ is non-decreasing and $\frac{f(x)}{1-x}$ is non-increasing.
We will prove that for $x<y$, $\{x,y\}\cap(1-\delta,1+\delta)\neq\emptyset$, and $1\not\in[x,y]$,
(A.12) $\exists\alpha_{\delta}>0\quad\textrm{s.t.}\quad\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}^{2}\geq\alpha_{\delta}(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)}.$
When $1-2\delta^{\prime}<x<y<1+2\delta^{\prime}$, (A.12) can be obtained with $\alpha_{\delta_{1}}=\frac{f^{\prime\prime}(1)(1-2\delta^{\prime})^{2}}{4}$ from the following two inequalities
(A.13) $f^{\prime}(y)-f^{\prime}(x)$ $=\int_{x}^{y}f^{\prime\prime}(\xi)\mathrm{d}\xi\geq\frac{f^{\prime\prime}(1)}{2}(y-x)$
[…]
$=f^{\prime\prime}(1)(y-x)$
$\leq 2(f^{\prime}(y)-f^{\prime}(x)).$
Here in the last inequality, we used (A.13).
When $1<x<1+\delta^{\prime}<1+2\delta^{\prime}<y$, (A.12) can be obtained with $\alpha_{\delta_{2}}=\frac{\delta^{\prime}f^{\prime\prime}(1)}{6}$ from the following two inequalities
(A.15) $f^{\prime}(y)-f^{\prime}(x)$ $=\int_{x}^{y}f^{\prime\prime}(\xi){\rm d}\xi\geq\int_{1+\delta^{\prime}}^{1+2\delta^{\prime}}f^{\prime\prime}(\xi)\mathrm{d}\xi\geq\frac{\delta^{\prime}f^{\prime\prime}(1)}{2}\geq\frac{\delta^{\prime}f^{\prime\prime}(1)}{2}\bigl{(}\frac{1}{x}-\frac{1}{y}\bigr{)},$
$\frac{f(x)}{1-x}-\frac{f(y)}{1-y}$ $=\int_{1}^{y}f^{\prime\prime}(\xi)\frac{y-\xi}{y-1}{\rm d}\xi-\int_{1}^{x}f^{\prime\prime}(\xi)\frac{x-\xi}{x-1}{\rm d}\xi$
(A.16) $=\frac{y-x}{(y-1)(x-1)}\int_{1}^{x}f^{\prime\prime}(\xi)(\xi-1){\rm d}\xi+\int_{x}^{y}f^{\prime\prime}(\xi)\frac{y-\xi}{y-1}{\rm d}\xi$
$\leq\delta^{\prime}f^{\prime\prime}(1)+\int_{x}^{y}f^{\prime\prime}(\xi){\rm d}\xi$
$\leq 3\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}.$
Here in the last inequality, we used (A.15).
When $10<x<1-2\delta^{\prime}<1-\delta<y<$ 1, (A.12) can be obtained with $\alpha_{\delta_{3}}=\frac{\alpha}{4}$. In fact, the right hand side of (A.12) can be bounded as
(A.17) $(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)}\leq\frac{f(x)}{x},$
where we take $y=1$ in the left hand side to obtain the above inequality. For the left hand side of eq. A.12, using (A.11) leads to
(A.18) $-f^{\prime}(x)$ $=\int_{x}^{1}f^{\prime\prime}(\xi){\rm d}\xi\geq\int_{1-2\delta^{\prime}}^{1}f^{\prime\prime}(\xi)\mathrm{d}\xi\geq\delta^{\prime}f^{\prime\prime}(1)\geq-2f^{\prime}(1-\delta)\geq-2f^{\prime}(y).$
Using Equations A.18, A.17 and A.8 leads to
Therefore, we can choose the $\delta>0$ defined in eqs. A.10 and A.11. For $x\in(1-\delta,1+\delta)$, $y>0$, $x\neq y$, and $1\not\in[x,y]$ we have (A.12) holds with $\alpha_{\delta}=\min\{\frac{f^{\prime\prime}(1)(1-2\delta^{\prime})}{4},\frac{\delta^{\prime}f^{\prime\prime}(1)}{6},\frac{\alpha}{4}\}$.
For $x\in(1-\delta,1+\delta)$ and $y>0$, the remaining cases are (i) $x=y$ and (ii) $1\in[x,y]$. For (i) (A.12) becomes $0\leq 0$; for (ii) (A.12) follows from 4.2. Hence, (4.4) holds with $\alpha_{\delta}>0$. The proof is complete. ∎
Proof of Lemma 4.3.

### v2

@@ 5.3 General results using the reverse $\chi^{2}$ divergence as the dual
(88) $\begin{split}&\int\Bigl(f^{\prime}\bigl(\frac{1}{\mu}\bigr)-f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)\Bigr)\Bigl(\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr)\rho\mathrm{d}\theta\\ \geq&\mu(1-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}f(\frac{1}{\mu})=\mu^{2}h(\frac{1}{\mu})\quad(\mu\geq 1+\delta_{f}\textrm{ and }h\textrm{ non-increasing},h\geq 0)\\ \geq&h\bigl(\frac{1}{1+\delta_{f}}\bigr)(\mu^{2}-1)=h(\frac{1}{1+\delta_{f}})\chi^{2}[{\rho^{*}}\|\rho].\end{split}$
According to the definition of $\delta_{f}$ in (84), we have
(89) $h\bigl(\frac{1}{1+\delta_{f}}\bigr)=\int^{1}_{\frac{1}{1+\delta_{f}}}f^{\prime\prime}(x)(1-x)\mathrm{d}x\geq\frac{f^{\prime\prime}(1)}{2}\int^{1}_{\frac{1}{1+\delta_{f}}}(1-x)\mathrm{d}x=\frac{\delta_{f}^{2}f^{\prime\prime}(1)}{4(1+\delta_{f})^{2}}.$
@@
In summary, combining the two scenarios, namely $\mu\leq 1+\delta_{f}$ in (86) and $\mu\geq 1+\delta_{f}$ in (LABEL:eq:generalized-gradient-dominant-condition-2b), we conclude that (79b) holds for
$\alpha_{\chi^{2},f}=\min\Bigl\{\alpha_{\chi^{2},f}^{\prime},\frac{\delta_{f}^{2}f^{\prime\prime}(1)}{4(1+\delta_{f})^{2}}\Bigr\}>0.$
This completes the proof.
5.4 Tighter results using the *-conjugate $f$-divergence as the dual
@@
In the last subsection, we obtain general convergence results by choosing $\bar{f}$ as the reverse $\chi^{2}$ divergence. The constant $\alpha_{f}$ depends on $f$ only. However, it depends on some implicit constants such as $\delta_{f},\alpha_{\chi^{2},f}$ that are dependent on $f$. The goal of this subsection is to improve the estimate for specific cases and obtain more explicit constants.
@@
We take ${\bar{f}}(x)=xf\bigl(\frac{1}{x}\bigr)$, known as as the *-conjugate function of $f$ [osterreicher2002csiszar]. For this choice, the $\bar{f}$-divergence has the following dual form
(90) $D_{{\bar{f}}}[\rho\|{\rho^{*}}]=\int{\rho^{*}}{\bar{f}}\Bigl(\frac{\rho}{{\rho^{*}}}\Bigr)\,\mathrm{d}\theta=\int\rho f\Bigl(\frac{{\rho^{*}}}{\rho}\Bigr)\,\mathrm{d}\theta.$
@@
Along the Fisher-Rao gradient flow (25) of $D_{f}$, we have
@@
The following theorem shows that the dual gradient dominance condition holds with an explicit constant when $f^{\prime\prime}$ is a polynomial.
@@
Given any convex function $f$ specified by $f^{\prime\prime}(x)=x^{p}$ with $f(1)=0$, when $p\leq-2$ or $1\leq p$, we have
(91) $-\int\Bigl(-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl[f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr]\Bigr)\Bigl(f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr)\mathrm{d}\theta\geq D_{\bar{f}}[\rho\|{\rho^{*}}]+D_{f}[\rho\|{\rho^{*}}],$
and when 2<p<1-2<p<-1, we have
(92) $-\int\Bigl(-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl[f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr]\Bigr)\Bigl(f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr)\mathrm{d}\theta\geq\frac{1}{2}\Bigl(D_{\bar{f}}[\rho\|{\rho^{*}}]+D_{f}[\rho\|{\rho^{*}}]\Bigr).$
@@
a consequence of Theorem 5.5, we get a convergence result for the Fisher-Rao gradient flow associated with the KL-divergence. More precisely, for such gradient flow, if we choose an initial density $\rho_{0}$ satisfying that ${\rm KL}[\rho_{0}\|{\rho^{*}}]+{\rm KL}[{\rho^{*}}\|\rho_{0}]<\infty$, then the Fisher-Rao gradient flow (assuming it is well-posed) converges exponentially fast with
@@
Proof 5.6 (Proof of Theorem 5.5).
First, we observe that both sides of these dual gradient dominance inequalities (91) and (92) remain the same under the transformation $f(x)\rightarrow f(x)+c(x-1)$ for any $c$. Therefore, for each $\rho$, it suffices to prove that (91) and (92) hold for one representative element under the transformation.
When p=1p=-1, we consider $f(x)=x\log x$; the corresponding $f$-divergence is the KL-divergence. We have the following
(93) $-\int\Bigl(-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl[f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr]\Bigr)\Bigl(f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr)\mathrm{d}\theta$
@@
Hence, (91) holds.
When p=2p=-2, we consider $f(x)=-\log x$; the corresponding $f$-divergence is the reverse KL-divergence. We have
(94) $-\int\Bigl(-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl[f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr]\Bigr)\Bigl(f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr)\mathrm{d}\theta$
@@
Hence, (91) holds.
When $p\neq-1$ and $p\neq-2$, we consider $f(x)=\frac{x^{p+2}-1}{(p+2)(p+1)}$. The corresponding $f$-divergences include $\chi^{2}$ divergence (when $p=0$). We have
(95) $-\int\Bigl(-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl[f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr]\Bigr)\Bigl(f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr)\mathrm{d}\theta$
@@
When p<2p<-2 or 1<p-1<p, applying Jensen’s inequality to the convex function $x^{p+2}~(x>0)$, we have
[…]
Therefore, (91) holds using the following estimate:
(96) $\frac{1}{(p+2)(p+1)}$ $\Bigl(\int\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}\mathrm{d}\theta\int\frac{{\rho^{*}}^{p+2}}{\rho^{p+1}}\mathrm{d}\theta-1\Bigr)$
@@
When 2<p<1-2<p<-1, applying Jensen’s inequality to the concave function $x^{p+2}~(x>0)$, we have
[…]
Note that $(p+1)(p+2)<0$. We get (92) using the following inequality
(97) $\frac{1}{(p+2)(p+1)}$ $\Bigl(\int\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}\mathrm{d}\theta\int\frac{{\rho^{*}}^{p+2}}{\rho^{p+1}}\mathrm{d}\theta-1\Bigr)$
@@
This concludes the proof.
6 Conclusion
@@
Furthermore, we have introduced the dual gradient dominance condition and demonstrated its validity for a broad class of $f$-divergences, notably including the KL divergence. The constants in this condition depend solely on $f$, independent of the target distribution $\rho^{*}$}. This characteristic sharply contrasts with the log-Sobolev inequality developed in the Wasserstein geometry, where constants depend on the target distribution.
@@
Future research directions include establishing the well-posedness of the Fisher-Rao gradient flows and exploring the properties of more general, including non-smooth, energy functionals within the Fisher-Rao geometry as well as the related Wasserstein-Fisher-Rao geometry. Designing efficient numerical approximations for these Fisher-Rao gradient flows, particularly for sampling applications as discussed in [lu2019accelerating, lu2022birth, huang2022efficient, lindsey2022ensemble, chen2024efficient, maurais2024sampling], remains a promising and open area of research.
@@
Proof A.1 (Proof of Lemma 4.5).